The money that we earn from the tickets is not supposed to to be more that 120% of 484013,92$.
Their are 2000 seats for the concert.These seats are divides in 3 categories:A,B,C.
In the categorie B,there are 600 more seats than in the categorie A.
The nomber of seats in the categorie C is the double of the nomber of seats in the categorie A.
The consert is for 7 days and new poeple come every day.
Each day,a ticket is ready to be sold for each seat.
We see that only 3/5 of the tickets for each categorie will be sold.
The minimum price of a ticket is 20$.
The price of a ticket in the categorie B is 15$ more than the ticket for the categorie C.
The ticket for the categorie B has to be equal to the half of a ticket of the categorie A.
How much money do we earn from the sold tickets?

Answer by rfer(16322)

2000 seats in 3 cats 3/5 sold for 7 days
A=x
B=x+600
C=2x
x+x+600+2x=2000
4x+600=2000
4x=1400
x=350 seats
A=x=350(.6)=210
B=x+600=950
C=2x=700
prices
a=35(2)=70
b=20+15=35
c=20
A=350(.6)=210(70)=$14700
B=950(.6)=570(35)=$19950
C=700(.6)=420(20)=$ 8400
total sales =$43050(7 days)=$301350